Present value with periodic

rates.

  Sam​ Hinds, a local​ dentist, is going to remodel the dental reception area and add two new workstations. He has contacted​ A-Dec, and the new equipment and cabinetry will cost

What will Sam have to pay for this equipment if the loan calls for

quarterlyquarterly

payments

​(44

per​ year)?

​$15 comma 00015,000.

The purchase will be financed with an interest rate of

8.58.5​%

loan over

77

years. What will Sam have to pay for this equipment if the loan calls for

quarterlyquarterly

payments

​(44

per​ year) and

weeklyweekly

payments

​(5252

per​ year)? Compare the annual cash outflows of the two payments. Why does the

weeklyweekly

payment plan have less total cash outflow each​ year?
Answer #1
QUARTERLY PAYMENTS
Pv Cost of new equipment $15,000
Rate Quarterly interest rate=(8.5/4)% 2.13%
Nper Number of payments=7*4 28
PMT Quarterly Payment $716.31 (Using PMT function with Rate=2.13%, Nper=28,Pv=-15000)
WEEKLY PAYMENTS
Pv Cost of new equipment $15,000
Rate Weekly interest rate=(8.5/52)% 0.16%
Nper Number of payments=7*52 364
PMT Quarterly Payment $54.71 (Using PMT function with Rate=0.16%, Nper=364,Pv=-15000)
Annual Cash Flow for quarterly payments $2,865.24 (716.31*4)
Annual Cash Flow for weekly payments $2,844.91 (54.71*52)
Weekly payment has lower Annual Cash Flow, because payments are made earlier than quarterly payments

Clipboard Font Alignment Number Styles : X V f =PMT(F6,F7,-F5) F G H I I K L QUARTERLY PAYMENTS Pv Rate Cost of new equipmentF15 =PMT(F13,F14,-F12) F G H I J QUARTERLY PAYMENTS Pv $15,000 2.13% Rate Nper PMT Cost of new equipment Quarterly interest r